OpenStudy (anonymous):

3) WILL GIVE MEDAL AND FAN!! Use the quadratic formula to solve x2 + 7x + 8 = 0. Estimate irrational solutions to the nearest tenth. a. {–12, 5} b. {–11.1, –2.9} c. {–5.6, –1.4} d. {–3.9, 12.1}

OpenStudy (wolf1728):

a = 1 b = 7 c = 8 x = [-7 +-sqroot (7*7 - 4*1*8)]/2*1 x = [-7 +-sqroot (17] / 2 x1 = [-7 +4.1231056256] / 2 x1 = -2.8768943744 / 2 x1 = -1.4384471872 x1 (rounded to nearest tenth) = -1.4 x2 = [-7 -4.1231056256] / 2 x2 = -11.1231056256 / 2 x2 = -5.5615528128 x2 (rounded to nearest tenth) = -5.6